NON-ADIABATIC ENERGIES OF THE HYDROGEN MOLECULE

\underline{KRZYSZTOF PACHUCKI}, Faculty of Physics, University of Warsaw, Hoża 69, 00-681 Warsaw, Poland; JACEK KOMASA, Faculty of Chemistry, A. Mickiewicz University, Grunwaldzka 6, 60-780 Poznań, Poland.

A novel perturbation theory has been developed to account for nonadiabatic effects in diatomic molecules^{a}. All molecular levels can be obtained simultanously from a single nuclear equation corrected for the presence of (m/M)^{2} terms. Comparison with the direct nonadiabatic calculations for rotationless states of H_{2} demonstrates an agreement at the level of 10^{−5} cm^{−1} for the ground tone transition.

aK. Pachucki and J. Komasa, J. Chem. Phys. 130, 164113 (2009)